OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool to the foot. A. 14 B. 8 C. 11 D. 10

OpenStudy (campbell_st):

\[206 = \pi r^2\] \[r = \sqrt{206/\pi}\] as an estimate.. 206/3 = 68 square root of 68 is approx 8...